OpenStudy (anonymous):

Simplify 10 over quantity of 8 minus 5 i

OpenStudy (nurali):

\[\frac{ 10 }{ 8-5i }\]

OpenStudy (anonymous):

\[\frac{ 10 }{8+5\iota }\] rationalise by multiplying the numerator and denominator by conjugate of denominator.

OpenStudy (nurali):

Multiply both side by 8+5i

OpenStudy (nurali):

\[\frac{ 10 }{ 8-5i }*\frac{ 8+5i }{ 8+5i }\]

OpenStudy (anonymous):

correction it is 8-5 iota

OpenStudy (anonymous):

thank you, guys (:

OpenStudy (nurali):

My Pleasure.